1. Determine a region of the xy-plane for which the given differential equation would have a unique solution whose graph passes through a point (x0, y0)in the region.

(4-y^2 ) y prime = x^2

A. A unique solution exists in the entire xy-plane.
B. A unique solution exists in the regions y < −2, −2 < y < 2, and y > 2.
C. A unique solution exists in the region consisting of all points in the xy-plane except (0, 2) and (0, −2).
D.A unique solution exists in the region y > −2.
E. A unique solution exists in the region y < 2.
1
Expert's answer
2020-06-17T19:10:07-0400

"(4-y^2)y'=x^2"

According to theorem of Existence of Unique Solution if "f(x,y)" and "df\/dy" are continious

on rectangular region "R" theh there is exist interval "I" on which unique exists.


We have:

"f(x,y)=\frac {x^2}{4-y^2}"

"\frac {df}{dy}=\frac {2yx^2}{(4-y^2)^2}"


"f(x,y)" and "df\/dy" are not continious at "y=-2" and "y=2"


Answer: C

A unique solution exists in the region consisting of all points in the xy-plane except "(0,2)" and "(0,-2)"